OpenStudy (anonymous):

Please, PLEASE help! Linda owes $3,861 on a credit card with a 22.3% interest rate compounded monthly. What is the monthly payment she should make in order to pay off this debt in 2 years, assuming she does not charge any more purchases with the card? I am not sure which formula to use ):

OpenStudy (anonymous):

Here are the answer choices: $160.88 $200.87 $361.93 $867.93

OpenStudy (the_fizicx99):

P(1+r)^x where x is time

OpenStudy (anonymous):

Thank you! Can you stay a bit to check my answer?

OpenStudy (the_fizicx99):

I tried pluging the numbers in and it gave me a big number, trying to see where I went wrong.

OpenStudy (anonymous):

Okay. I honestly don't remember this formula though

OpenStudy (the_fizicx99):

No wrong formula, its um: interest = 0.223 annual interest rate. time = 0.223/12 = 0.0185833... monthly rate. rate = 1+j = 1.0185833

OpenStudy (anonymous):

Now that I am reading it over again, would I used the monthly payments formula?

OpenStudy (the_fizicx99):

Yes

OpenStudy (the_fizicx99):

do you understand the process used?

OpenStudy (anonymous):

Yes, I do. Hold on and I'll show you what I ended up with

OpenStudy (the_fizicx99):

Here: 3861^(24)(1-k)/(1-k^(24)), k=1+.223/12

OpenStudy (the_fizicx99):

I just used k, i.e: K = 1 + R/ 12

OpenStudy (anonymous):

Oh okay. I just don't recognize the formula. I thought the monthly payments formula was M=B [ i (1+ i)^nt/ (1 + i)^nt - 1)]

OpenStudy (the_fizicx99):

Yes, the PMT formula :D

OpenStudy (anonymous):

is that what you have trying to tell me? haha

OpenStudy (the_fizicx99):

Yes, now can you solve? I can't do it for you, I've already been given a warning for doing so..

OpenStudy (anonymous):

Yes, and this is what I got: 0.0520264224085132

OpenStudy (anonymous):

It's nothing close to any of the answer choices :l

OpenStudy (anonymous):

Do I have to do something with the 3,861 with the answer I got above?

OpenStudy (the_fizicx99):

0.0520264224085132 * 3861 = 200.8740169192694652 *cough* get it?

OpenStudy (anonymous):

OH! Okay! I always mess that last step up! I kept adding it! lol

OpenStudy (anonymous):

Thank you very much!

OpenStudy (the_fizicx99):

Yw, what grade are you in?

OpenStudy (anonymous):

12 grade

OpenStudy (the_fizicx99):

9th. Learning Algebra II on the side. Sorry for confusing you at some point

OpenStudy (anonymous):

This is Advanced Algebra with Financial Applications. I'd say you're doing pretty well for a freshmen haha Wish I could understand math like that lol

OpenStudy (the_fizicx99):

I should've skipped a grade when I had the chance, I'd be a junior :/

OpenStudy (anonymous):

I may need more help...mind if I ask for your help again if needed? and I didn't know you could still skip a grade

OpenStudy (the_fizicx99):

Um I'm kinda doing hw on the side too lol xD

OpenStudy (the_fizicx99):

Make a new question and tag a mod or ambi and they'll help you :p GL!!

OpenStudy (anonymous):

I'm sorry! I didn't know lol That's okay, I'm sure I can find some else. (: Good luck with that homework!

OpenStudy (the_fizicx99):

Yw, make new post, this is laggy af. Earth Science mixed with physics smh

OpenStudy (anonymous):

okay lol